OpenStudy (anonymous):

Let F(x)=f(x^ 3 ) and G(x)=(f(x)) ^3 . You also know that a ^2 =4,f(a)=2,f ′ (a)=2,f ′ (a ^3 )=6 . Find F'(a) and G'(a)

OpenStudy (anonymous):

\[G(x)=(f(x))^3\] \[G'(x)=3f(x)^2\times f'(x)\]

OpenStudy (anonymous):

\[G'(a)=3f(a)^2\times f'(a)=3\times 2^2\times 2\]